# Unconventional criticality in $O(D)$-invariant loop-constrained Landau theory

### v2

@@ Abstract
Then, the two-point function to the order $b^{2}$ in momentum space,
@@
Here, $\Delta_{ij}^{\left(0\right)}\left(\bm{r}\right)$ is the free polarization propagator in real space, see Appendix LABEL:app:PropReal, and, for an isotropic system,
Although the diagram $\mathcal{A}_{kl}$ (14) representing the amputated correlation function is not transverse, the full second-order propagator (13) is transverse as required by the ${\mbox{$\nabla$}}\cdot\bm{P}=0$ constraint. Performing the integral in (14) and substituting the diagram into (13) yields for $D=3$,
@@
where $C_{1}=11/6$, $C_{2}=\left(792\ln 2-315\ln 3-181\right)/540$, and $c$ contains the UV cutoff of the real-space integrals, i.e. $c\to 0$. Note that in the massless theory ($a=0$) the diagram (14) is logarithmically divergent for $D=3$, hindering the calculation of $\eta$. This issue does not arise in massive theory.
Next, we rewrite the propagator (16) in the form
where $\eta$ is calculated at the fixed point. Differentiation with respect to $\Lambda$ is performed keeping $b$, but not $a$, fixed. More precisely, the process or renormalization involves the replacement of bare quantities by renormalized ones, which are overall independent of $\Lambda$ and are given as a perturbative expansion in terms of the bare parameters of the theory. The requirement that all renormalized correlation functions are $\Lambda$-independent yields the RG functions [35]. Explicitly, we have,
It remains to be seen whether the universality class uncovered here can be connected to some fractionalization mechanism, by representing $\bm{P}$ as a composite operator, similarly to the CP1 representation in quantum antiferromagnets [1]. DQC is indeed facilitated by a mechanism where staggered Berry phases interfere with quadrupled instanton events, which ultimately causes the order parameter to fall apart into spinons [1, 4].
Conclusions and final remarks – In this Letter, we have shown that imposing a local divergence-free constraint on the polarization field drives a ferroelectric system into a universality class that lies beyond the conventional LGW framework. This profoundly modifies the configuration space, restricting it to loop-like polarization structures and thereby altering the critical behavior. The RG analysis demonstrates that the polarization correlation function features a large anomalous dimension in three dimensions, $\eta=0.239$, which significantly exceeds that of ordinary $O(3)$ LGW models.
This unconventional scaling originates from an emergent $U(1)$ gauge symmetry, which becomes explicit when the constraint is resolved as $\bm{P}={\mbox{$\nabla$}}\times\bm{A}$, rendering the effective theory manifestly gauge invariant. In this sense, the zero divergence constraint is understood as a local conservation law corresponding to a local gauge symmetry. This leads to the emergence of a gauge field and therefore to a large value of $\eta$. Usually, the argument runs the opposite way, with a gauge symmetry yielding a local conservation law, one example being current conservation in electrodynamics. From a field theoretic perspective, at the critical point current conservation implies that currents do not have an anomalous dimension [40]. This fact is a consequence of the scale invariance of the critical point. The opposite is true here: the order parameter, viewed as a conserved quantity, has a large anomalous dimension.
Acknowledgements
Appendix D: Tadpole correction to the propagator – The one-loop propagator is given by,
@@ Acknowledgements
$=$ $\frac{1}{q^{2}+a}\left(\delta_{in}-\frac{q_{i}q_{n}}{q^{2}}\right)$
[…]
where $\int_{k}\equiv\int\frac{d^{D}k}{(2\pi)^{D}}$ and
Appendix E: Free polarization proparator in real space – The constraint ${\mbox{$\nabla$}}\cdot\bm{P}=0$ imposes that in the absence of interactions the polarization propagator is given in momentum space by Eq. (11). To obtain the corresponding propagator in real space, we rewrite the propagator as,
and introduce the function,
$C(r;a)$ $=$ $\int\frac{d^{D}q}{(2\pi)^{D}}\frac{e^{i\bm{q}\cdot\bm{r}}}{q^{2}+a}$ (30)
$=$ $\frac{1}{(2\pi)^{D/2}}\left(\frac{\sqrt{a}}{r}\right)^{(D-2)/2}K_{\frac{D-2}{2}}(\sqrt{a}r),$
which for $D=3$ holds exactly. Then, in real space,
$=$ $\frac{1}{4\pi ar^{D}}\left(\frac{1}{2\pi}\right)^{\frac{D-3}{2}}$
[…]
For $D=3$, this expression simplifies to
allowing us to calculate the two-loop propagator (16).

### v3

@@ Acknowledgements
Appendix F: Two-loop propagator – The two-point function to the order $b^{2}$ in momentum space,
@@
Here, $\Delta_{ij}^{\left(0\right)}\left(\bm{r}\right)$ is the free polarization propagator in real space, see Appendix LABEL:app:PropReal. Note that in the massless theory ($a=0$) the diagram (35) is logarithmically divergent for $D=3$. This issue does not arise in massive theory.
Although the diagram $\mathcal{A}_{kl}$ (35) representing the amputated correlation function is not transverse, the full second-order propagator (34) is transverse as required by the constraint. Performing the integral in (35) and substituting into (34) yields for $D=3$ the expression (14) with $C_{1}=11/6$, $C_{2}=\left(792\ln 2-315\ln 3-181\right)/540$.
Appendix G: Explicit derivation of the anomalous dimension – The first nonvanishing contribution to the field renormalization constant, $Z$, is given by the coefficient of the $q^{2}$ term in the renormalized two-point vertex function corresponding to the two-loop propagator (14,15) as
@@
$\quad Z\left(a,b\right)=-\frac{1}{\tilde{\Delta}^{2}}\frac{\partial{\tilde{\Delta}}}{\partial{q^{2}}}\bigg|_{q^{2}=0}=1+\frac{C_{2}}{4\pi^{2}}\frac{b^{2}}{a}.$ (36)
The RG function $\eta\left(a,b\right)$ is given by [26],
$\eta\left(a,b\right)=-\Lambda\frac{\partial}{\partial{\Lambda}}\bigg|_{b_{R}{\text{ fixed}}}{\ln Z}\simeq-\frac{1}{Z}\frac{C_{2}}{4\pi^{2}}\frac{2b^{2}}{a}.$ (37)
Note that the differentiation with respect to $\Lambda$ is performed keeping the (renormalized) coupling constant, but not the mass, fixed. Specifically, we replace the bare quantities $a,b$ by $\Lambda$-independent renormalized ones $a_{R},b_{R}$, which are then given as a perturbative expansion in terms of the bare parameters of the theory. Next, recalling the dimensionless coupling constants, $\hat{a}=\Lambda^{-2}a$ and $\hat{b}=N_{D}\Lambda^{D-4}b$, gives $b^{2}/a=N_{D}^{-2}\Lambda^{2(3-D)}\hat{b}^{2}/\hat{a}$, or, in three dimensions, $b^{2}/a=N_{3}^{-2}\hat{b}^{2}/\hat{a}=4\pi^{4}\hat{b}^{2}/\hat{a}$. The values of the dimensionless coupling constants at the fixed point, $\hat{a}_{*}$ and $\hat{b}_{*}$, are obtained from the RG equations (9). We solve $\Lambda d\hat{b}/d\Lambda=0$ to get $\hat{b}_{*}$ and then tune the mass parameter $\hat{a}$ to criticality using the RG $\beta$-function for $\hat{a}$. This yields
$\hat{a}_{*}=\frac{\left(D-4\right)\left(D+2\right)^{2}}{D^{3}+2\left(D+2\right)^{2}};\,\,\hat{b}_{*}=\frac{2D\left(4-D\right)\left(D+2\right)}{\left(D-1\right)\left(D^{2}+10D+12\right)},$ (38)
resulting in $\hat{a}_{*}=-25/77$ and $\hat{b}_{*}=5/17$ for $D=3$. Substituting the fixed point values to the RG function (37) leads to the expression (16), and, eventually, to the numerical value $\eta\approx 0.239$.